OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (ghazi):

see the line is parallel to the y=-x+6 therefore it will have the same slope as of the given line and equation of the new line would be y=-x+D now to find D substitute (3,4) at x and y you'll get D

OpenStudy (anonymous):

x + y = 7?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

Thank you for the help?

OpenStudy (ghazi):

anytime